# Universal mechanistic rules for de novo design of enzymes

### v1

@@ III.1 Preliminary considerations
It is crucial to note that the components of the mobility tensor $M_{\alpha\beta}$ are determined by the geometric features of the enzyme and molecule, which can be tuned as to optimize the function of the enzyme. The basic principle is illustrated in Fig. 2(a), for the example of dissociation of a dimer molecule. When a catalytic reaction takes place (the internal phase $\phi_{\textrm{e}}$ is kicked over the chemical energy barrier by thermal noise), the coupling to the molecule length arising from the conformational changes of the enzyme drives a corresponding evolution of the molecule length, as represented by $\phi_{\textrm{m}}$, and in particular can cause $\phi_{\textrm{m}}$ to overcome its energy barrier. The evolution along the two-dimensional energy landscape $V_{\textrm{e}}(\phi_{\textrm{e}})+V_{\textrm{m}}(\phi_{\textrm{m}})$ is shown in Fig. 2(b), where the black path corresponds to a catalyzed reaction.
We expect to have a strong coupling between the fuel-to-waste ($\phi_{\textrm{e}}$) and substrate-to-product ($\phi_{\textrm{m}}$) reactions when the off-diagonal components in the mobility tensor are non-negligible. In the limiting cases where either the middle sub-unit is large compared to the enzyme-only sub-unit, $\mu_{\textrm{b}}/\mu_{1}\rightarrow 0$, or the conformational changes of the enzyme are small compared to the dissociation length of the molecule, $\ell_{\textrm{e}}/\ell_{\textrm{m}}\rightarrow 0$, the off-diagonal terms vanish, leading to a decoupling of the dynamics between the enzyme and the molecule. To achieve a strong coupling, we would aim to maximize $\mu_{\textrm{b}}/\mu_{1}$ and $\ell_{\textrm{e}}/\ell_{\textrm{m}}$, from which one can infer the first two golden rules [(i) and (ii)] for designing an enzyme discussed above. In the following, we define the ratio $h\equiv\mu_{\textrm{b}}/\mu_{1}=\mu_{\textrm{b}}/(\mu_{\textrm{e}}+\mu_{\textrm{b}})$ to be the “coupling strength”, and we note that it is bounded as $0\leq h\leq 1$. For $h<0.5$, the mobility of the shared sub-unit is larger than that of the enzyme-only sub-unit, and the reverse is true for $h>0.5$.
@@ III.2 Deterministic dynamics and global bifurcation
Before considering the full stochastic (noise-activated) dynamics of the system, we first study the deterministic dynamics given by Eq. (1), as we can apply tools from dynamical systems theory to obtain an intuitive understanding of the expected behaviour, particularly in the experimentally-relevant case of low noise, where $k_{\rm B}T$ is smaller than all other energy scales.
@@
In Fig. 3, phase portraits of the deterministic dynamics are shown for low and high coupling strengths and both types of reaction (see Appendix C for details of the numerical calculation of phase portraits). Depending on the type of reaction, different sets of fixed points appear in the dynamics (stable fixed points are depicted as circles, saddle points as diamonds, and unstable fixed points as asterisks). For the dissociation reaction, there are four fixed points (one stable, two saddle points, one unstable) since there is no local minimum of the molecule potential after the energy barrier. The stable fixed point at $\phi_{\textrm{m}}=0$ corresponds to the substrate (dimer) state, whereas $\phi_{\textrm{m}}\to\infty$ corresponds to the product (dissociated) state. For the switch reaction, there are six fixed points (two stable, three saddle points, one unstable). The stable fixed point at $\phi_{\textrm{m}}=-1$ corresponds to the substrate (short) state, whereas that at $\phi_{\textrm{m}}=+1$ corresponds to the product (long) state. It is worth noting that the positions of the fixed points do not depend on the coupling, but only on the potentials $V_{\textrm{e}}(\phi_{\textrm{e}})$ and $V_{\textrm{m}}(\phi_{\textrm{m}})$. Thus, changes in the geometry of the enzyme-molecule complex (and thus in the coupling strength) leave the location and nature of the fixed points unchanged. As a consequence, there cannot be any local bifurcations in the dynamical system and only global bifurcations are allowed [35]. By considering the topology of the different basins of attraction in the dynamical system, we note that global bifurcations occur between Figs. 3(a) and 3(b) for the dissociation reaction, and between Figs. 3(c) and 3(d) for the switch reaction, respectively.
@@
The phase portraits allow us to uncover the different possibilities emerging in the full stochastic dynamics by classifying the different types of possible deterministic trajectories, depicted as solid flow lines in Fig. 3, and considering the likelihood of their stochastic activation by random thermal kicks, depicted as dashed flow lines in Fig. 3, which will take the system from an initial state around a stable fixed point to the domain of interest. Trajectories starting in the cyan region in Fig. 3 are attracted to the fixed point at $(\phi_{\textrm{e}},\phi_{\textrm{m}})=(0,0)$ (for the dissociation reaction) or $(0,-1)$ (for the switch reaction). On the other hand, those starting in the green region are attracted to the fixed point at $(\phi_{\textrm{e}},\phi_{\textrm{m}})=(2\pi,0)$ (dissociation) or $(2\pi,-1)$ (switch). This class of trajectories, which are depicted as solid white flow lines in Fig. 3, represent the process in which the molecule remains in the substrate state and the enzyme performs a futile cycle. Analogously, trajectories starting in the magenta region end at $(0,\infty)$ (dissociation) or $(0,1)$ (switch), corresponding to spontaneous transition of the molecule to the product state, without an accompanying transition in the enzyme, while those in the orange region converge to $(2\pi,\infty)$ (dissociation) or $(2\pi,1)$ (switch), representing a process in which the molecular transition coincides with the enzymatic cycle. The relative frequency of the activation of the different processes corresponding to these distinct classes will depend on the likelihood of the availability of the thermal kick of the right strength that is needed to initiate each of them.
@@
Figure 4: Critical coupling strength at which the bifurcation occurs, as a function of (a) the geometric ratio $\ell_{\textrm{e}}/\ell_{\textrm{m}}$ and the ratio $E_{*}/E_{\mathrm{ba}}$ describing the internal enzyme reaction, and (b) the mobility ratio $\mu_{1}/\mu_{2}$ and $E_{*}/E_{\mathrm{ba}}$. In the white region outside the black solid line, a bifurcation does not occur within the physical range $0\leq h\leq 1$.
@@
Figure 5: Marginal steady-state distributions for the switching reaction, (a,c) at low noise $k_{\rm B}T/\Delta=0.15$, and (b,d) at high noise $k_{\rm B}T/\Delta=0.4$. The distribution for the enzyme coordinate $\phi_{\textrm{e}}$ is shown in (a,b), and that for the molecule coordinate $\phi_{\textrm{m}}$ in (c,d). In (a,b) the black line labelled “NESS” is the analytical result for an uncoupled enzyme given in Eq. (D). In (c,d) the black line labelled “Boltzmann” is the equilibrium distribution $P_{\textrm{m}}(\phi_{\textrm{m}})\propto\exp(-V_{\textrm{m}}/k_{\rm B}T)$.
@@
More specifically, let us consider the case of weak coupling in Figs. 3(a,c). Thermal fluctuations can most likely kick the system over one of the two saddle points or transition states, either the one to the right of the stable fixed point shown as red diamond (white trajectory) or the one above it shown as grey diamond (yellow trajectory). When traversing the red diamond, the enzyme will complete an internal reaction ending at $\phi_{\textrm{e}}=2\pi$, while the molecule remains in the substrate state. When traversing the grey diamond, the molecule undergoes a reaction and turns into product, while the enzyme completes no internal reaction and remains at $\phi_{\textrm{e}}=0$. Thus, the internal (fuel-to-waste) reaction of the enzyme and the molecule (substrate-to-product) reaction are completely uncoupled in these cases. Less likely thermal fluctuations can kick the system across the unstable fixed point shown as asterisk (blue trajectory), leading to a simultaneous occurrence of both the enzyme and the molecule cycles.
@@
We next consider the case of strong coupling, after the global bifurcation has occurred and the basins of attraction have rearranged; see Fig. 3(b,d) and Supplemental Videos 5 and 6. Thermal fluctuations kicking the system over the grey diamond (yellow trajectory) will still result in molecular reactions without associated internal reactions in the enzyme. However, if the system is activated over the red diamond, the more likely scenario will be for the enzyme to complete an internal reaction simultaneously with the molecular reaction (blue trajectory) unlike in the weak coupling case. The two reactions thus become coupled, and in particular, a fuel-to-waste reaction inside the enzyme now triggers a substrate-to-product reaction for the molecule. The futile enzyme cycle (white trajectory) is the more unlikely scenario in this case. We note that a similar bifurcation occurs for the reverse switching process, where the energetically favourable and initial state for the molecule is the long conformation; see Appendix B.
@@
The occurrence of the bifurcation is crucial for the coupling between the two processes to emerge, and thus for the enzyme to become active. We thus explore how the critical coupling depends on the geometric and energetic parameters of the enzyme-substrate complex. In Fig 4, we show the critical value of the coupling $h$ as a function of the ratio $E_{*}/E_{\mathrm{ba}}$ of energies in the fuel-to-waste reaction, and $\ell_{\textrm{e}}/\ell_{\textrm{m}}$ or $\mu_{1}/\mu_{2}$ which are directly related to the geometrical properties of the complex. In Fig 4(a), it is observed that the bifurcation takes place at low coupling strengths when the driving $E_{*}$ is large, which is linked to higher reaction rates [related to golden rule (iii)]. In addition, there is a non-trivial behaviour as a function of the length ratio $\ell_{\textrm{e}}/\ell_{\textrm{m}}$ at a fixed value of $E_{*}/E_{\mathrm{ba}}$, showing non-monotonic behaviour with a minimum at intermediate values. In Fig 4(b), we see that the bifurcation is favoured at higher values of the ratio $\mu_{1}/\mu_{2}$. The results in Fig. 4 serve to highlight the different parameters that must be tuned in order to favour a bifurcation in the deterministic dynamics. Below, we show how these results translate to the full noise-activated stochastic dynamics.
@@ III.3.1 Steady state: driving a reaction uphill
Passive enzymes can lower reaction barriers, but do not shift reaction equilibria, which are still purely governed by the free energy differences between substrate and product. Fuelled enzymes, on the other hand, can use the free energy provided by the fuel to drive the substrate-to-product reaction uphill, i.e. favouring the formation of a high free energy product from a low free energy substrate. To study whether our model enzyme is capable of driving reactions uphill, we focus on the switch reaction (Fig. 1(d)), and numerically solve the Fokker-Planck equation (Eq. (5), see Appendix C for details of the numerical solution), where we are interested in how the probability distribution evolves over time and in the form of the steady state distribution in the long time limit, where $\partial_{t}P_{\mathrm{ss}}\rightarrow 0$.
@@
In stark contrast, at low noise the coupling $h$ has a substantial effect on the NESS attained by the molecule, $P_{\textrm{m}}(\phi_{\textrm{m}})$; see Fig. 5(c). At $h=0$, the distribution $P_{\textrm{m}}(\phi_{\textrm{m}})$ corresponds to a Boltzmann distribution $\propto e^{-V_{\textrm{m}}(\phi_{\textrm{m}})/k_{\rm B}T}$, and thus shows two peaks (corresponding to the short and long states), with the peak corresponding to the lower energy short state being much higher. As the coupling $h$ increases, we see this trend progressively reverse, as the fuelled action of the enzyme drives more and more probability towards the long state. This demonstrates that the fuelled action of the enzyme can indeed drive a reaction uphill, in the thermodynamically unfavourable direction. However, the action of the enzyme becomes much less effective at high noise, where deviations from the Boltzmann equilibrium are minimal; see Fig. 5(d).
@@
where $\phi_{\mathrm{m},\mathrm{max}}$ is the location of the energy barrier of the molecular reaction along the $\phi_{\textrm{m}}$ coordinate. We consider an enzyme and a molecule that are initially separated from each other, and bind at time $t=0$. We thus initialize the enzyme probability distribution as being in the steady state of an uncoupled enzyme ($h=0$). For the molecule, we consider two cases: (i) the molecule starts in its Boltzmann equilibrium, or (ii) the molecule starts only in the short state (with probability distributed in a narrow Gaussian around it). Results from numerical solution of the Fokker-Planck equation (Eq. (5)) are shown in Fig. 6. For $h=0$, $P_{\mathrm{long}}$ tends to the probability corresponding to the equilibrium Boltzmann distribution, as expected. However, for $h>0$, $P_{\mathrm{long}}$ is consistently higher than what the Boltzmann distribution dictates. Clearly, coupling to the fuelled enzyme can drive the molecular reaction uphill. The effect is most pronounced at larger coupling strengths and lower noise.
@@
The behaviour of $P_{\mathrm{long}}(t)$ in Fig. 6 can be well approximated by an exponential relaxation $P_{\mathrm{long}}(t)\simeq P_{\mathrm{long}}(\infty)-[P_{\mathrm{long}}(\infty)-P_{\mathrm{long}}(0)]\exp(-k_{\mathrm{cat}}t)$, where $k_{\mathrm{cat}}$ is the inverse timescale associated with the transient regime in Fig. 6. With this choice, Eq. (13) evaluates to
$\langle P_{\mathrm{long}}\rangle\simeq\frac{k_{\mathrm{off}}P_{\mathrm{long}}(0)+k_{\mathrm{cat}}P_{\mathrm{long}}(\infty)}{k_{\mathrm{off}}+k_{\mathrm{cat}}}.$ (14)
In the limiting case in which $k_{\mathrm{off}}\ll k_{\mathrm{cat}}$, so that the typical timescale for unbinding is longer than the timescale of the transient regime in Fig. 6, we find $\langle P_{\mathrm{long}}\rangle\simeq P_{\mathrm{long}}(\infty)$.
@@
and is constant in the NESS. The rate of energy dissipation is then given by $\Omega_{\mathrm{e}}E_{*}$. In Fig. 7, we show the steady state values of $P_{\mathrm{long}}$ and the energy dissipation rate $\tilde{\Omega}_{\mathrm{e}}E_{*}/k_{\rm B}T$ (where $\tilde{\Omega}_{\mathrm{e}}$ is the rate in dimensionless time units $\tilde{t}$ as introduced above), as a function of the coupling $h$ and the geometric parameter $\ell_{\textrm{e}}/\ell_{\textrm{m}}$. Interestingly, there are large regions of parameter space where the enzyme is highly functional (larger $P_{\mathrm{long}}$) while the energy dissipation is much lower than the corresponding value for an uncoupled enzyme, which can be explained by a slow-down of the fuelled enzymatic reaction (smaller $\Omega_{\textrm{e}}$) due to the coupling to the molecular reaction. It is also worth noting that $P_{\mathrm{long}}$ is non-monotonic as a function of $\ell_{\textrm{e}}/\ell_{\textrm{m}}$, and peaks at intermediate values of order one, i.e. when the deformation of the enzyme is comparable to the deformation required of the molecule during the course of the reaction.

### v2

@@ III.1 Preliminary considerations
It is crucial to note that the components of the mobility tensor $M_{\alpha\beta}$ are determined by the geometric features of the enzyme and molecule, which can be tuned as to optimize the function of the enzyme. The basic principle is illustrated in Figure 2(A), for the example of dissociation of a dimer molecule. When a fuel-to-waste reaction takes place (the internal phase $\phi_{\textrm{e}}$ is kicked over the chemical energy barrier by thermal noise), the coupling to the molecule length arising from the conformational changes of the enzyme drives a corresponding evolution of the molecule length, as represented by $\phi_{\textrm{m}}$, and in particular can cause $\phi_{\textrm{m}}$ to overcome its energy barrier. The evolution along the two-dimensional energy landscape $V_{\textrm{e}}(\phi_{\textrm{e}})+V_{\textrm{m}}(\phi_{\textrm{m}})$ is shown in Figure 2(B), where the black path corresponds to a catalyzed reaction.
We expect to have a strong coupling between the fuel-to-waste ($\phi_{\textrm{e}}$) and substrate-to-product ($\phi_{\textrm{m}}$) reactions when the off-diagonal components in the mobility tensor are non-negligible. In the limiting cases where either the middle sub-unit is large compared to the enzyme-only sub-unit, $\mu_{\textrm{b}}/\mu_{1}\rightarrow 0$, or the conformational changes of the enzyme are small compared to the dissociation length of the molecule, $\ell_{\textrm{e}}/\ell_{\textrm{m}}\rightarrow 0$, the off-diagonal terms vanish, leading to a decoupling of the dynamics between the enzyme and the molecule. To achieve a strong coupling, we would aim to maximize $\mu_{\textrm{b}}/\mu_{1}$ and $\ell_{\textrm{e}}/\ell_{\textrm{m}}$, from which one can infer the first two golden rules [(i) and (ii)] for designing an enzyme discussed above. In the following, we define the ratio $h\equiv\mu_{\textrm{b}}/\mu_{1}=\mu_{\textrm{b}}/(\mu_{\textrm{e}}+\mu_{\textrm{b}})$ to be the “coupling strength”, and we note that it is bounded as $0\leq h\leq 1$. For $h>0.5$, the mobility of the shared sub-unit is larger than that of the enzyme-only sub-unit, and the reverse is true for $h<0.5$. To achieve a fast conformational change and thus, infer the rule (iii), we would aim to maximize the chemical driving force of the fuel-to-waste process $E_{*}$ or increase the mobility $\mu_{\phi}$ governing this process.
@@ III.2 Deterministic dynamics and global bifurcation
Before considering the full stochastic (noise-activated) dynamics of the system, we first study the deterministic dynamics given by Equation (1), as we can apply tools from dynamical systems theory to obtain an intuitive understanding of the expected behaviour, particularly in the experimentally-relevant case of low noise, where $k_{\rm B}T$ is smaller than all other energy scales.
@@
In Figure 3, phase portraits of the deterministic dynamics are shown for low and high coupling strengths and both types of reaction (see Methods for details of the numerical calculation of phase portraits). Depending on the type of reaction, different sets of fixed points appear in the dynamics (stable fixed points are depicted as circles, saddle points as diamonds, and unstable fixed points as asterisks). For the dissociation reaction, there are four fixed points (one stable, two saddle points, one unstable) since there is no local minimum of the molecule potential after the energy barrier. The stable fixed point at $\phi_{\textrm{m}}=0$ corresponds to the substrate (dimer) state, whereas $\phi_{\textrm{m}}\to\infty$ corresponds to the product (dissociated) state. For the switch reaction, there are six fixed points (two stable, three saddle points, one unstable). The stable fixed point at $\phi_{\textrm{m}}=-1$ corresponds to the substrate (short) state, whereas that at $\phi_{\textrm{m}}=+1$ corresponds to the product (long) state. It is worth noting that the positions of the fixed points do not depend on the coupling, but only on the potentials $V_{\textrm{e}}(\phi_{\textrm{e}})$ and $V_{\textrm{m}}(\phi_{\textrm{m}})$. Thus, changes in the geometry of the enzyme-molecule complex (and thus in the coupling strength) leave the location and nature of the fixed points unchanged. As a consequence, there cannot be any local bifurcations in the dynamical system and only global bifurcations are allowed [36]. By considering the topology of the different basins of attraction in the dynamical system, we note that global bifurcations occur between Figures 3(A) and 3(B) for the dissociation reaction, and between Figures 3(C) and 3(D) for the switch reaction, respectively.
@@
The phase portraits allow us to uncover the different possibilities emerging in the full stochastic dynamics by classifying the different types of possible deterministic trajectories, depicted as solid flow lines in Figure 3, and considering the likelihood of their stochastic activation by random thermal kicks, depicted as dashed flow lines in Figure 3, which will take the system from an initial state around a stable fixed point to the domain of interest. Trajectories starting in the cyan region in Figure 3 are attracted to the fixed point at $(\phi_{\textrm{e}},\phi_{\textrm{m}})=(0,0)$ (for the dissociation reaction) or $(0,-1)$ (for the switch reaction). On the other hand, those starting in the green region are attracted to the fixed point at $(\phi_{\textrm{e}},\phi_{\textrm{m}})=(2\pi,0)$ (dissociation) or $(2\pi,-1)$ (switch). This class of trajectories, which are depicted as solid white flow lines in Figure 3, represent the process in which the molecule remains in the substrate state and the enzyme performs a futile cycle. Analogously, trajectories starting in the purple region end at $(0,\infty)$ (dissociation) or $(0,1)$ (switch), corresponding to spontaneous transition of the molecule to the product state, without an accompanying transition in the enzyme, while those in the gold region converge to $(2\pi,\infty)$ (dissociation) or $(2\pi,1)$ (switch), representing a process in which the molecular transition coincides with the enzymatic cycle. The relative frequency of the activation of the different processes corresponding to these distinct classes will depend on the likelihood of the availability of the thermal kick of the right strength that is needed to initiate each of them.
@@
Figure 4: Critical coupling strength at which the bifurcation occurs (A) as a function of the geometric ratio $\ell_{\textrm{e}}/\ell_{\textrm{m}}$ and the ratio $E_{*}/E_{\mathrm{ba}}$ describing the internal enzyme reaction, and (B) the mobility ratio $\mu_{1}/\mu_{2}$ and $E_{*}/E_{\mathrm{ba}}$. In the white region outside the black solid line, a bifurcation does not occur within the physical range $0\leq h\leq 1$.
@@
Figure 5: Marginal steady-state distributions for the switching reaction (A,C) at low noise $k_{\rm B}T/\Delta=0.15$, and (B,D) at high noise $k_{\rm B}T/\Delta=0.4$. The distribution for the enzyme coordinate $\phi_{\textrm{e}}$ is shown in (A,B), and that for the molecule coordinate $\phi_{\textrm{m}}$ in (C,D). In (A,B) the black line labelled “NESS” is the analytical result for an uncoupled enzyme given in Equation (S9). In (C,D) the black line labelled “Boltzmann” is the equilibrium distribution $P_{\textrm{m}}(\phi_{\textrm{m}})\propto\exp(-V_{\textrm{m}}/k_{\rm B}T)$.
@@
More specifically, let us consider the case of weak coupling in Figure 3(A,C). Thermal fluctuations can most likely kick the system over one of the two saddle points or transition states, either the one to the right of the stable fixed point shown as red diamond (white trajectory) or the one above it shown as grey diamond (yellow trajectory). When traversing the red diamond, the enzyme will complete an internal reaction ending at $\phi_{\textrm{e}}=2\pi$, while the molecule remains in the substrate state. When traversing the grey diamond, the molecule undergoes a reaction and turns into product, while the enzyme completes no internal reaction and remains at $\phi_{\textrm{e}}=0$. Thus, the internal (fuel-to-waste) reaction of the enzyme and the molecule (substrate-to-product) reaction are completely uncoupled in these cases. Less likely thermal fluctuations can kick the system across the unstable fixed point shown as an asterisk (blue trajectory), leading to a simultaneous occurrence of both the enzyme and the molecule cycles. In this weak coupling regime, traversing either of the saddle points (red or grey diamonds) leads to effectively one-dimensional dynamics [see Figure 3(A, C)] since only one of the two coordinates advances. Thus, dynamics in this regime are analogous to a Kramers-type one-dimensional barrier-crossing process.
@@
We next consider the case of strong coupling, after the global bifurcation has occurred and the basins of attraction have rearranged; see Figure 3(B,D) and Supplemental Videos 5 and 6. Thermal fluctuations kicking the system over the grey diamond (yellow trajectory) will still result in molecular reactions without associated internal reactions in the enzyme. However, if the system is activated over the red diamond, the more likely scenario will be for the enzyme to complete an internal reaction simultaneously with the molecular reaction (blue trajectory) unlike in the weak coupling case. The two reactions thus become coupled, and in particular, a fuel-to-waste reaction inside the enzyme now triggers a substrate-to-product reaction for the molecule. The futile enzyme cycle (white trajectory) is the more unlikely scenario in this case. Importantly, the dynamics become two-dimensional since the two reaction coordinates become strongly coupled and influence each other. This occurs even if the underlying energy landscape remains unchanged, through the emergence of cross-mobility terms in the dynamics of the reaction coordinates. Thus, the present dynamical systems theory approach is crucial in understanding catalysis in this system. We note that a similar bifurcation occurs for the reverse switching process, where the energetically favourable and initial state for the molecule is the long conformation; see Supplemental Information and Figure S1.
@@
The occurrence of the bifurcation is crucial for the coupling between the two processes to emerge, and thus for the enzyme to become active. We thus explore how the critical coupling depends on the geometric and energetic parameters of the enzyme-substrate complex. In Fig 4, we show the critical value of the coupling $h$ as a function of the ratio $E_{*}/E_{\mathrm{ba}}$ of energies in the fuel-to-waste reaction, and $\ell_{\textrm{e}}/\ell_{\textrm{m}}$ or $\mu_{1}/\mu_{2}$ which are directly related to the geometrical properties of the complex. In Fig 4, it is observed that the bifurcation takes place at low coupling strengths when the driving $E_{*}$ is large, which is linked to higher reaction coordinate speed $\dot{\phi}_{\textrm{e}}$ and therefore a faster conformational change of the enzyme, in connection with the golden rule (iii). In addition, there is a non-trivial behaviour as a function of the length ratio $\ell_{\textrm{e}}/\ell_{\textrm{m}}$ at a fixed value of $E_{*}/E_{\mathrm{ba}}$, showing non-monotonic behaviour with a minimum at intermediate values. In Fig 4(B), we see that the bifurcation is favoured at higher values of the ratio $\mu_{1}/\mu_{2}$. The results in Figure 4 serve to highlight the different parameters that must be tuned in order to favour a bifurcation in the deterministic dynamics. Below, we show how these results translate to the full noise-activated stochastic dynamics.
@@ III.3.1 Steady state: driving a reaction uphill
Passive enzymes can lower reaction barriers, but do not shift reaction equilibria, which are still purely governed by the free energy differences between substrate and product. Fuelled enzymes, on the other hand, can use the free energy provided by the fuel to drive the substrate-to-product reaction uphill, i.e. favouring the formation of a high free energy product from a low free energy substrate. To study whether our model enzyme is capable of driving reactions uphill, we focus on the switch reaction (Figure 1(E)), and numerically solve the Fokker-Planck equation (Equation (5), see Methods for details of the numerical solution), where we are interested in how the probability distribution evolves over time and in the form of the steady state distribution in the long time limit, where $\partial_{t}P_{\mathrm{ss}}\rightarrow 0$.
@@
In stark contrast, at low noise the coupling $h$ has a substantial effect on the NESS attained by the molecule, $P_{\textrm{m}}(\phi_{\textrm{m}})$; see Figure 5(C). At $h=0$, the distribution $P_{\textrm{m}}(\phi_{\textrm{m}})$ corresponds to a Boltzmann distribution $\propto e^{-V_{\textrm{m}}(\phi_{\textrm{m}})/k_{\rm B}T}$, and thus shows two peaks (corresponding to the short and long states), with the peak corresponding to the lower energy short state being much higher. As the coupling $h$ increases, we see this trend progressively reverse, as the fuelled action of the enzyme drives more and more probability towards the long state. This demonstrates that the fuelled action of the enzyme can indeed drive a reaction uphill, in the thermodynamically unfavourable direction. However, the action of the enzyme becomes much less effective at high noise, where deviations from the Boltzmann equilibrium are minimal; see Figure 5(D).
@@
where $\phi_{\mathrm{m},\mathrm{max}}$ is the location of the energy barrier of the molecular reaction along the $\phi_{\textrm{m}}$ coordinate. We consider an enzyme and a molecule that are initially separated from each other, and bind at time $t=0$. We thus initialize the enzyme probability distribution as being in the steady state of an uncoupled enzyme ($h=0$). For the molecule, we consider two cases: (i) the molecule starts in its Boltzmann equilibrium, or (ii) the molecule starts only in the short state (with probability distributed in a narrow Gaussian around it). Results from numerical solution of the Fokker-Planck equation (Equation (5)) are shown in Figure 6. For $h=0$, $P_{\mathrm{long}}$ tends to the probability corresponding to the equilibrium Boltzmann distribution, as expected. However, for $h>0$, $P_{\mathrm{long}}$ is consistently higher than what the Boltzmann distribution dictates. Clearly, coupling to the fuelled enzyme can drive the molecular reaction uphill. The effect is most pronounced at larger coupling strengths and lower noise. Interestingly, $P_{\mathrm{long}}$ can change nonmonotonically with $h$, as seen in Figure 6(B), where $h=0.67$ shows a smaller $P_{\mathrm{long}}$ than $h=0.48$. This is most clearly seen in Figure 7(A), where we show the steady state values of $P_{\mathrm{long}}$ as a function of $h$ and the geometric ratio $\ell_{\mathrm{e}}/\ell_{\mathrm{m}}$. The nonmonotonicity as a function of these two quantities can be understood from the nonlinear form in which these quantities appear in the mobilities $M_{\mathrm{ee}}$ and $M_{\mathrm{me}}$, which moreover determine both the deterministic and stochastic behavior of the system. Overall, Figure 7(A) serves as a validation of the golden rules (i) and (ii).
@@
The behaviour of $P_{\mathrm{long}}(t)$ in Figure 6 can be well approximated by an exponential relaxation $P_{\mathrm{long}}(t)\simeq P_{\mathrm{long}}(\infty)-[P_{\mathrm{long}}(\infty)-P_{\mathrm{long}}(0)]\exp(-t/\tau_{\mathrm{ss}})$, where $\tau_{\mathrm{ss}}$ is the timescale associated with the transient regime in Figure 6. With this choice, Equation (13) evaluates to
$\langle P_{\mathrm{long}}\rangle\simeq\frac{k_{\mathrm{off}}P_{\mathrm{long}}(0)+P_{\mathrm{long}}(\infty)/\tau_{\mathrm{ss}}}{k_{\mathrm{off}}+1/\tau_{\mathrm{ss}}}.$ (14)
In the limiting case in which $k_{\mathrm{off}}\tau_{\mathrm{ss}}\ll 1$, so that the typical timescale for unbinding is longer than the timescale of the transient regime in Figure 6, we find $\langle P_{\mathrm{long}}\rangle\simeq P_{\mathrm{long}}(\infty)$.
@@
and is constant in the NESS. The rate of energy dissipation is then given by $\Omega_{\mathrm{e}}E_{*}$. In Figure 7(B), we show the energy dissipation rate $\tilde{\Omega}_{\mathrm{e}}E_{*}/k_{\rm B}T$ (where $\tilde{\Omega}_{\mathrm{e}}$ is the rate in dimensionless time units $\tilde{t}$ as introduced above), as a function of the coupling $h$ and the geometric parameter $\ell_{\textrm{e}}/\ell_{\textrm{m}}$. Interestingly, there are large regions of parameter space where the enzyme is highly functional (larger $P_{\mathrm{long}}$) while the energy dissipation is much lower than the corresponding value for an uncoupled enzyme, which can be explained by a slow-down of the fuelled enzymatic reaction (smaller $\Omega_{\textrm{e}}$) due to the coupling to the molecular reaction.
Figure 7: Results on steady-state distributions and energy dissipation (A) The normalized probability of finding the shape-switching molecule in the long state, where $P_{\mathrm{long}}^{\mathrm{eq}}\simeq 0.383$ is the probability at equilibrium of finding the shape-switching molecule in the product state. (B) Energy dissipation rate, versus the length ratio $\ell_{\textrm{e}}/\ell_{\textrm{m}}$ and the coupling strength $h$. The noise strength is $k_{\rm B}T/\Delta=0.2$.